# SVL: Spike-based Vision-language Pretraining for Efficient 3D Open-world Understanding

### v1

@@ 4.3 Re-parameterizable Vision-language Integration
After pretraining, the text prompt is applied to guide the learned spike-based representation to enable the following zero-shot transfer tasks. However, during inference, the traditional methods [14; 15] still require a high-parameter text encoder, which compromises the spike-driven nature and significantly impacts the deployment of neuromorphic hardware. To address this issue, we propose a novel structure called the re-parameterizable Vision-language module (Rep-VLI), which converts text embeddings into the weights of linear layers. It can be described as:
$W^{L}_{i}=e^{\tau}\mathcal{E}_{\theta}^{T}({T}_{i}^{t}),$ (9)
where $\mathcal{E}_{\theta}^{T}(\cdot)$ is the text encoder and $W_{i}$ is the re-parameterized layer weights on the $i$-th input text prompt ${T}_{i}^{t}$ at $t$ time step. Then we concatenate the re-parameterized weights $W_{i}$ by $K$ text prompts to $W^{L}\in\mathbb{R}^{K\times C}$. And the classification logits are calculated with the spike trains $\mathcal{F}^{S}$ and re-parameterized weights as:
$\text{logits}_{i}=\text{softmax}[\frac{W^{L}}{T}\mathcal{E}^{S}_{\theta}({D}^{t}_{i})],$ (10)
where $\mathcal{E}_{\theta}^{S}(\cdot)$ is the spike-based encoder, $T$ is the time step, and ${D}^{t}_{i}$ is the input 3D data. Rep-VLI overcomes the limitations of traditional high-parameter models by removing the need for a text encoder during inference. It re-parameterizes text into a lightweight format aligned with the spike-driven framework, ensuring efficient deployment on neuromorphic hardware while supporting diverse tasks like 3D open-world detection and segmentation [36; 37] with spike-based encoders.
@@ 5 Experiments
To validate that our spike-based 3D encoder learns robust visual representations via SVL, we evaluate it on diverse 3D open-world tasks, including zero-shot classification and visual question answering. The pretrained encoder is also fine-tunable for downstream tasks like 3D classification, segmentation, detection, and action recognition. This section details the experimental setup, including backbones, datasets, and implementation, followed by quantitative results and ablation studies on modality count, time steps, and loss functions. For further details about implementation and the datasets, please refer to Appendix E and Appendix F, respectively.
3D Backbone Networks
We evaluate three 3D backbones within our SVL framework: (i) Spike PointNet [26], a lightweight, energy-efficient spiking model for point clouds; (ii) E-3DSNN [5], a state-of-the-art model leveraging spike sparse convolution for sparse 3D inputs; and (iii) Spike PointFormer is our proposed Transformer-based SNN backbone for 3D data. Further details are provided in Appendix B. We use OpenCLIP’s [38] ViT-G/14 text encoder (frozen during pretraining) to define a shared feature space, into which the 3D modality is aligned. During inference, our Rep-VLI transforms text embeddings into compact weights, preserving the spike-driven nature of the above spike-based encoders.
Architecture Model Pre-train Method Input $T\times D$ Point+Text Param (M) Energy (mJ) Obj. M40. ANN PointCLIP [19] N/A Image N/A 25.5+57.3 24.9+20.3 1.9 20.2 PointNet [22] Openshape [16] Point N/A 3.47+202.5 20.1+71.7 24.4 74.9 Point-Bert [39] Point N/A 21.9+202.5 83.2+71.7 43.2 82.8 Sparseconv [40] Voxel N/A 5.3+202.5 0.61+71.7 31.7 78.8 Voxel N/A 41.3+202.5 2.13+71.7 43.4 83.4 Point-Bert [39] Ulip [14] Point N/A 21.9+227.8 81.2+80.6 34.9 69.6 Point-Bert [39] Ulip2 [15] Point N/A 21.9+202.5 83.7+71.7 50.6 84.7 SNN $\text{SpikeCLIP}^{*}$ [11] N/A Image 4$\times$1 9.5+22.8 10.6+0.41 0.5 5.1 Spike PointNet [25] SVL (Ours) Point 1$\times$4 3.57 0.27 24.9 76.3 Spike PointFormer Point 1$\times$4 22.1 9.4 43.4 83.1 E-3DSNN-T [5] Voxel 1$\times$4 2.1 0.04 33.6 79.6 E-3DSNN-S [5] Voxel 1$\times$4 3.5 0.09 36.4 81.3 E-3DSNN-L [5] Voxel 1$\times$4 17.7 0.64 43.9 84.6 E-3DSNN-H [5] Voxel 1$\times$4 46.7 0.79 47.0 85.4
@@
Table 1: 3D Zero-shot classification results on the large-scale Objaverse-LVIS (Obj.) [41] and ModelNet40 (M40.) [24] datasets. "*" denotes self-implementation results with open-source code. "Energy" denotes the estimated energy consumption, following [5; 42]; further details are provided in Appendix D. "Point+Text" denotes the parameters of the point encoder and the text encoder.
@@
We evaluate the zero-shot classification performance of our models on the widely-used ModelNet40 [24] and the larger, more challenging Objaverse-LVIS [41]. Compared to other benchmarks, Objaverse-LVIS offers broader class coverage and a long-tailed distribution, providing a more realistic evaluation of open-world 3D understanding [16]. As shown in Tab. 1, our SVL-based E-3DSNN achieves 85.4% accuracy on ModelNet40 with only 17.7M parameters, outperforming both ANN and SNN baselines. This demonstrates SVL’s effectiveness in enhancing both accuracy and efficiency.
Specifically, compared to OpenShape and ULIP, our model achieves 85.4% accuracy (vs. 83.4% and 69.6%), consumes only 0.79 mJ of energy (vs. 161.8 mJ and 73.8 mJ), and uses fewer parameters (17.7M vs. 41.3M and 21.9M). It also delivers a 15.8% accuracy gain over ULIP-based Point-BERT [14] while consuming just 11.4% of the energy. On Objaverse-LVIS, our model performs comparably to ULIP-2 [15] but with a 204$\times$ energy efficiency advantage. This is enabled by our Rep-VLI module, which reparameterizes text embeddings into compact, spike-compatible weights for zero-shot inference, preserving the spike-driven nature of the encoder. Compared to prior SNN approaches such as SpikeCLIP [11], SVL substantially improves the visual representation capacity of spike-based encoders in zero-shot 3D tasks.
Method Vision Encoder LLM Input S.-BERT SimCSE B-1. R-L. MET. InstructBLIP-13B [43] ViT [33] Vicua [44] Image 45.90 48.86 4.65 8.85 13.23 LLaVA-13B [45] Image 46.37 45.90 4.02 8.15 12.58 PointLLM-13B [12] PointBert [39] Vicua [44] Point 47.91 49.12 3.83 7.23 12.26 SVL-13B (Ours) Spike PointFormer Point 44.87 45.91 3.77 6.85 12.25 SVL-13B (Ours)* Spike PointFormer Point 47.80 47.08 11.45 14.69 16.40 Human N/A N/A N/A 100.00 100.00 100.00 100.00 100.00
Table 2: 3D object captioning results on Objaverse-LVIS. "*" indicates SVL-13B is prompted for shorter captions with no more than 20 words. The evaluation utilizes a range of metrics, including Sentence-BERT, SimCSE, BLEU-1, ROUGE-L, and METEOR.
@@
We combine the SVL-trained Spike PointFormer with a language model [44] via the LLaVA framework [45] for multimodal pre-training and fine-tuning (see Appendix B). On the 3D object captioning benchmark, prompted with “Describe this 3D model in detail,” our SVL-13B achieves performance comparable to state-of-the-art ANN methods (Tab. 2). Semantic metrics (Sentence-BERT, SimCSE) confirm strong alignment with human references. Notably, as the first SNN-based method in 3D captioning, SVL-13B achieves comparable annotation quality compared to PointLLM. We also evaluate on 3D question answering. As shown in Fig. 2, the model effectively interprets shape, material, function, and context, including visual and functional cues, while demonstrating commonsense reasoning. Despite lacking dense textures, SVL-13B achieves strong perception-language alignment, comparable to ANN models across diverse object types.
@@
We first fine-tuned our models on 3D classification datasets such as ModelNet40 [24] and ScanObjectNN [46] to evaluate the 3D visual representation capabilities acquired through SVL pretraining. As shown in Tab. 3, the SVL pretraining significantly enhances performance, with the E-3DSNN [5] and the Spike PointNet [25] architecture achieving improvements of 1.9% and 1.0%, respectively, on ModelNet40. On the more challenging ScanObjectNN dataset, the Spike PointNet architecture demonstrates a substantial accuracy increase, rising from 70.0% to 76.1%. Subsequently, we extended our fine-tuning experiments to datasets such as Semantic KITTI [47] and KITTI [48]. As illustrated in Tab. 2, our SVL pretraining delivers marked improvements in both 3D segmentation and detection tasks, with the E-3DSNN [5] exhibiting performance gains of 1.1% and 1.2%, respectively.
@@
We further fine-tune our SVL-pretrained spike-based encoder on DVS datasets, including DVS128 Gesture [49] and DVS Action [50], to assess spatiotemporal feature extraction. During pretraining, the I-LIF time step was set to 1 for efficiency, then increased to 6 during evaluation to better capture temporal dynamics. We adopt the point-based method from [23] for efficient DVS data processing (see Section 4.1). As shown in Tab. 4, SVL-pretrained E-3DSNN and Spike Point improve by 2.1% and 1.6% on DVS Action and DVS128 Gesture, respectively, indicating strong scalability and temporal modeling ability of SVL-trained SNNs.
Architecture Model Input Param (M) Energy (mJ) $T\times D$ ModelNet40 ScanObjectNN ANN PointNet [22] Point 3.27 2.02 N/A 89.2 68.2 PointNet + ULIP [14] Point 3.47 2.34 N/A 92.1 72.1 Pointformer [51] Point 4.91 5.1 N/A 92.8 81.3 SNN P2SResLNet [26] Point 14.3 - $4\times 1$ 88.7 81.2 SpikingPointNet [52] Point 3.47 0.91 $16\times 1$ 88.6 66.6 Spike PointNet [25] Point 3.47 0.24 $1\times 4$ 88.2 70.0 Spike PointNet + SVL Point 3.47 0.27 $1\times 4$ 90.1 ($\uparrow 1.9$) 76.1 ($\uparrow 6.1$) E-3DSNN-S [5] Voxel 3.27 0.02 $1\times 4$ 91.7 78.7 E-3DSNN-S + SVL Voxel 3.27 0.02 $1\times 4$ 92.7 ($\uparrow 1.0$) 80.9 ($\uparrow 2.2$) E-3DSNN-L [5] Voxel 17.7 0.26 $1\times 4$ 91.2 80.2 E-3DSNN-L + SVL Voxel 17.7 0.31 $1\times 4$ 93.7 ($\uparrow 2.5$) 83.0 ($\uparrow 2.8$)
Table 3: 3D Downstream Tasks: 3D classification results on ModelNet40 (M-40) [24] and ScanObjectNN (Scan-O) [46].
Architecture Method $T\times D$ KITTI AP-E (%) Semantic KITTI mIoU (%) DVS Action Acc. (%) DVS128 Gesture Acc. (%) ANN E-3DANN [5] N/A 89.4 69.4 - - PointNet [22] N/A - 14.6 75.1 95.3 SNN E-3DSNN [5] $1\times 4$ 89.6 68.5 - - E-3DSNN + SVL $1\times 4$ 90.7 ($\uparrow 1.1$) 69.7 ($\uparrow 1.2$) - - Spike PointNet [25] $1\times 4$ / $6\times 4$ 12.1 78.4 96.9 Spike PointNet + SVL $1\times 4$ / $6\times 4$ 15.6 ($\uparrow 2.1$) 80.5 ($\uparrow 2.1$) 98.5 ($\uparrow 1.6$)
Table 4: 3D Downstream Tasks: 3D segmentation, detection, and human action recognition results on KITTI [48], Semantic KITTI [47], DVS Action [50], and DVS128 Gesture [49]. Moreover, for the DVS dataset, we adopt a pre-training timestep of $1\times 4$, consistent with other datasets, and a fine-tuning timestep of $6\times 4$.

### v2

@@ 4.3 Re-parameterizable Vision-Language Integration
While pretrained vision-language models excel at zero-shot transfer, their inference phase typically requires a large, computationally expensive text encoder. This presents a significant bottleneck for SNNs, undermining their inherent efficiency and complicating deployment on neuromorphic hardware.
To overcome this, we introduce the Re-parameterizable Vision-Language Integration (Rep-VLI) module. Rep-VLI’s core innovation is to pre-compute and embed textual information directly into the weights of a lightweight classification layer, completely discarding the text encoder during inference.
Specifically, for a set of $K$ candidate text prompts $\{T_{1},T_{2},\dots,T_{K}\}$, we use the text encoder $\mathcal{E}_{\theta}^{T}$ to generate a corresponding weight matrix $W^{L}\in\mathbb{R}^{K\times C}$:
$W^{L}_{j}=\tau\mathcal{E}_{\theta}^{T}({T}_{j}),$ (10)
where $W^{L}_{j}$ is the re-parameterized weight vector for the $j$-th text prompt.
During inference, we adopt a hardware-friendly spike-count decision rule instead of a conventional softmax. For an input 3D data point $D_{i}$, our spike-based encoder $\mathcal{E}_{\theta}^{S}$ produces spike features $\mathbf{s}[t]\in\{0,1\}^{C}$ over $T$ timesteps. The predicted class, $\hat{y}_{i}$, is the one whose corresponding weight vector aligns best with the accumulated spike activity:
$\text{logits}_{i}=\arg\max_{j}\frac{1}{T}\sum_{t=1}^{T}W^{L}_{j}\cdot\mathcal{E}^{S}_{\theta}({D}^{t}_{i}),$ (11)
This process selects the class with the highest firing rate, preserving a fully spike-driven and hardware-compatible inference path. Ultimately, Rep-VLI elegantly sidesteps the need for a persistent text encoder at inference time. By re-parameterizing textual knowledge into a compact weight matrix, it ensures our model remains lightweight and perfectly aligned with the operational principles of spike-based neuromorphic systems.
4.4 Spike-driven 3D Encoder
Backbone Suite
We instantiate SVL with three spike-driven 3D backbones: (i) Spike PointNet (Wu et al., 2024) for lightweight point-cloud processing; (ii) E-3DSNN (Qiu et al., 2025a), a sparse spike-convolutional backbone suited to edge deployment; and (iii) Spike-driven PointFormer, our fully spike-driven Transformer for high-capacity cloud settings.
Spike-driven PointFormer
As shown in Fig. 3, given a point set $\mathcal{P}_{k}\in\mathbb{R}^{T\times N\times 3}$, we form local neighborhoods by farthest-point sampling and $k$-NN grouping:
$X=\mathrm{KNN}(\mathrm{FPS}(P)),\qquad X\in\mathbb{R}^{T\times N^{\prime}\times 3},$ (12)
A learnable add-only pointwise embedding (addition only since inputs are spikes), followed by an I-LIF neuron $\mathcal{SN}(\cdot)$, produces spike features:
$S=\mathcal{SN}\!\big(\mathrm{MLP}(X)\big),\qquad S\in\mathbb{R}^{T\times N^{\prime}\times D},$ (13)
We then perform local-to-global feature extraction with $L$ Spike-driven PointFormer layers ($\mathrm{SDF}$):
$f_{0}$ $=\mathrm{EMP}(S),$ (14)
$f_{\ell}$ $=\mathrm{SDF}(f_{\ell-1})+f_{\ell-1},\quad\ell=1,\dots,L,$
where $\mathrm{EMP}$ denotes element-wise max pooling over the neighborhood axis and all $f_{\ell}\in\mathbb{R}^{T\times N^{\prime}\times D}$.
Spike-driven self-attention inside $\mathrm{SDf}$. From $f_{\ell-1}$, three add-only linear maps yield $Q,K,V\in\mathbb{R}^{T\times N^{\prime}\times D}$, which are converted to spikes:
$Q_{S}=\mathcal{SN}(Q),\quad K_{S}=\mathcal{SN}(K),\quad V_{S}=\mathcal{SN}(V),$ (15)
A matrix-multiplication variant (3D-SDSA) used in Spike-driven PointFormer is
$\mathrm{SDSA}_{\text{3D}}(Q_{S},K_{S},V_{S})$ $=\mathcal{SN}\!\big(Q_{S}\,(K_{S}^{\top}V_{S})\big)$ (16)
$=\mathcal{SN}\!\big((Q_{S}K_{S}^{\top})\,V_{S}\big),$
Here all products are spike-driven matmuls that reduce to sparse additions via address-event accumulation algorithms (Horowitz, 2014), preserving end-to-end spike computation.
@@ 5 Experiments
Table 1: 3D Zero-shot classification results on the large-scale Objaverse-LVIS (Obj.) (Deitke et al., 2023) and ModelNet40 (M40.) (Wu et al., 2015) datasets. ”*” denotes self-implementation results with open-source code. ”Energy” denotes the estimated energy consumption, following (Qiu et al., 2025a; Yao et al., 2023); further details are provided in Appendix C. ”Point+Text” denotes the parameters of the point encoder and the text encoder.
[…]
Table 2: 3D object captioning results on Objaverse-LVIS. ”*” indicates SVL-13B is prompted for shorter captions with no more than 20 words. The evaluation utilizes a range of metrics, including Sentence-BERT, SimCSE, BLEU-1, ROUGE-L, and METEOR.
Method Vision Encoder LLM Input S.-BERT SimCSE B-1. R-L. MET. InstructBLIP-13B (Dai et al., 2023) ViT (Dosovitskiy, 2020) Vicua (Chiang et al., 2023) Image 45.90 48.86 4.65 8.85 13.23 LLaVA-13B (Liu et al., 2023a) Image 46.37 45.90 4.02 8.15 12.58 PointLLM-13B (Xu et al., 2024) PointBert (Yu et al., 2022) Vicua (Chiang et al., 2023) Point 47.91 49.12 3.83 7.23 12.26 PointLLM-13B* (Xu et al., 2024) PointBert (Yu et al., 2022) Point 50.15 50.83 17.09 20.99 16.45 SVL-13B (Ours) Spike-driven PointFormer-L Point 44.87 45.91 3.77 6.85 12.25 SVL-13B (Ours)* Spike-driven PointFormer-L Point 47.80 47.08 11.45 14.69 16.40 SVL-13B (Ours)* Spike-driven PointFormer-L SpikeLLM (Xing et al., 2025) Point 51.21 50.18 18.45 21.32 18.40 Human N/A N/A N/A 100.00 100.00 100.00 100.00 100.00
To validate that our spike-based 3D encoder learns robust visual representations via SVL, we evaluate it on diverse 3D open-world tasks, including zero-shot classification and visual question answering. The pretrained encoder is also fine-tunable for downstream tasks like 3D classification, segmentation, detection, and action recognition. This section details the experimental setup, including backbones, datasets, and implementation, followed by quantitative results and ablation studies on modality count, time steps, and loss functions. For additional speed and accuracy comparisons between our Spike-driven PointFormer and other spike-based 3D encoders—as well as temporal scene understanding results and CLIP encoder sizes—see Appendix A. Implementation details and dataset descriptions are provided in Appendix D and Appendix F, respectively.
Table 3: 3D Downstream Tasks: 3D classification results on ModelNet40 (M-40) (Wu et al., 2015) and ScanObjectNN (Scan-O) (Uy et al., 2019).
Architecture Model Input Param (M) Energy (mJ) $T\times D$ ModelNet40 ScanObjectNN ANN PointNet (Qi et al., 2017) Point 3.27 2.02 N/A 89.2 68.2 PointNet + ULIP (Xue et al., 2023) Point 3.47 2.34 N/A 92.1 72.1 Pointformer (Zhao et al., 2021) Point 4.91 30.1 N/A 92.8 81.3 SNN Spike Point TransFormer (Wu et al., 2025b) Point 9.6 21.1 4 $\times$1 88.5 80.1 P2SResLNet (Wu et al., 2024) Point 14.3 - $4\times 1$ 88.7 81.2 SpikingPointNet (Lan et al., 2023) Point 3.47 0.91 $16\times 1$ 88.6 66.6 Spike PointNet (Ren et al., 2024) Point 3.47 0.24 $1\times 4$ 88.2 70.0 Spike PointNet + SVL Point 3.47 0.27 $1\times 4$ 90.1 ($\uparrow 1.9$) 76.1 ($\uparrow 6.1$) E-3DSNN-S (Qiu et al., 2025a) Voxel 3.27 0.02 $1\times 4$ 91.7 78.7 E-3DSNN-S + SVL Voxel 3.27 0.02 $1\times 4$ 92.7 ($\uparrow 1.0$) 80.9 ($\uparrow 2.2$) E-3DSNN-L (Qiu et al., 2025a) Voxel 17.7 0.26 $1\times 4$ 91.2 80.2 E-3DSNN-L + SVL Voxel 17.7 0.31 $1\times 4$ 93.7 ($\uparrow 2.5$) 83.0 ($\uparrow 2.8$) Spike-driven PointFormer-S (Ours) Point 7.69 5.1 1 $\times$4 92.6 82.1 Spike-driven PointFormer-L (Ours) Point 22.1 9.4 1 $\times$4 92.1 81.7 Spike-driven PointFormer-L (ours) +SVL Point 22.1 9.8 1 $\times$4 93.9 ($\uparrow 1.8$) 83.4 ($\uparrow 1.7$)
Table 4: 3D Downstream Tasks: 3D segmentation, detection, and human action recognition results on KITTI (Geiger et al., 2012b), Semantic KITTI (Behley et al., 2019), DVS Action (Miao et al., 2019), and DVS128 Gesture (Amir et al., 2017). Moreover, for the DVS dataset, we adopt a pre-training timestep of $1\times 4$, consistent with other datasets, and a fine-tuning timestep of $6\times 4$.
Architecture Method $T\times D$ KITTI AP-E (%) Semantic KITTI mIoU (%) DVS Action Acc. (%) DVS128 Gesture Acc. (%) ANN E-3DANN (Qiu et al., 2025a) N/A 89.4 69.4 - - PointNet (Qi et al., 2017) N/A - 14.6 75.1 95.3 SNN E-3DSNN (Qiu et al., 2025a) $1\times 4$ 89.6 68.5 - - E-3DSNN + SVL $1\times 4$ 90.7 ($\uparrow 1.1$) 69.7 ($\uparrow 1.2$) - - Spike PointNet (Ren et al., 2024) $1\times 4$ / $6\times 4$ - 12.1 78.4 96.9 Spike PointNet + SVL $1\times 4$ / $6\times 4$ - 15.6 ($\uparrow 2.1$) 80.5 ($\uparrow 2.1$) 98.5 ($\uparrow 1.6$)
@@
We evaluate the zero-shot classification performance of our models on the widely-used ModelNet40 (Wu et al., 2015) and the larger, more challenging Objaverse-LVIS (Deitke et al., 2023). Compared to other benchmarks, Objaverse-LVIS offers broader class coverage and a long-tailed distribution, providing a more realistic evaluation of open-world 3D understanding (Liu et al., 2023b). As shown in Tab. 1, our SVL-based E-3DSNN achieves 85.4% accuracy on ModelNet40 with only 17.7M parameters, outperforming both ANN and SNN baselines. This demonstrates SVL’s effectiveness in enhancing both accuracy and efficiency.
[…]
@@
We combine the SVL-trained Spike PointFormer with a language model (Chiang et al., 2023) via the LLaVA framework (Liu et al., 2023a) for multimodal pre-training and fine-tuning (see Appendix H). On the 3D object captioning benchmark, prompted with “Describe this 3D model in detail,” our SVL-13B achieves performance comparable to state-of-the-art ANN methods (Tab. 2). Semantic metrics (Sentence-BERT, SimCSE) confirm strong alignment with human references. Notably, as the first SNN-based method in 3D captioning, SVL-13B achieves comparable annotation quality compared to PointLLM. We also evaluate on 3D question answering. As shown in Fig. 2, the model effectively interprets shape, material, function, and context, including visual and functional cues, while demonstrating commonsense reasoning. Despite lacking dense textures, SVL-13B achieves strong perception-language alignment, comparable to ANN models across diverse object types.
@@
We first fine-tuned our models on 3D classification datasets such as ModelNet40 (Wu et al., 2015) and ScanObjectNN (Uy et al., 2019) to evaluate the 3D visual representation capabilities acquired through SVL pretraining. As shown in Tab. 5, the SVL pretraining significantly enhances performance, with the E-3DSNN (Qiu et al., 2025a) and the Spike PointNet (Ren et al., 2024) architecture achieving improvements of 1.9% and 1.0%, respectively, on ModelNet40. On the more challenging ScanObjectNN dataset, the Spike PointNet architecture demonstrates a substantial accuracy increase, rising from 70.0% to 76.1%. Subsequently, we extended our fine-tuning experiments to datasets such as Semantic KITTI (Behley et al., 2019) and KITTI (Geiger et al., 2012b). As illustrated in Tab. 2, our SVL pretraining delivers marked improvements in both 3D segmentation and detection tasks, with the E-3DSNN (Qiu et al., 2025a) exhibiting performance gains of 1.1% and 1.2%, respectively.
@@
We further fine-tune our SVL-pretrained spike-based encoder on DVS datasets, including DVS128 Gesture (Amir et al., 2017) and DVS Action (Miao et al., 2019), to assess spatiotemporal feature extraction. During pretraining, the I-LIF time step was set to 1 for efficiency, then increased to 6 during evaluation to better capture temporal dynamics. We adopt the point-based method from (Wang et al., 2019) for efficient DVS data processing (see Section 4.1). As shown in Tab. 4, SVL-pretrained E-3DSNN and Spike Point improve by 2.1% and 1.6% on DVS Action and DVS128 Gesture, respectively, indicating strong scalability and temporal modeling ability of SVL-trained SNNs.
Comparisons between different Backbone
Table 5: 3D Downstream Tasks: 3D classification results on ModelNet40 (M-40) (Wu et al., 2015) and ScanObjectNN (Scan-O) (Uy et al., 2019).
Architecture Model Input Param (M) Energy (mJ) $T\times D$ ModelNet40 ScanObjectNN ANN PointNet (Qi et al., 2017) Point 3.27 2.02 N/A 89.2 68.2 Pointformer (Zhao et al., 2021) Point 4.91 30.1 N/A 92.8 81.3 Spike Point TransFormer (Wu et al., 2025b) Point 9.6 21.1 4 $\times$1 88.5 80.1 KPConv (Thomas et al., 2019) Point 14.3 - N/A 92.9 85.3 3DShapeNets (Wu et al., 2015) Voxel 6.97 0.61 N/A 88.2 - E-3DANN-S (Qiu et al., 2025a) Voxel 3.27 0.13 $1\times 4$ 91.7 79.7 SNN P2SResLNet (Wu et al., 2024) Point 14.3 - $4\times 1$ 88.7 81.2 SpikingPointNet (Lan et al., 2023) Point 3.47 0.91 $16\times 1$ 88.6 66.6 Spike PointNet (Ren et al., 2024) Point 3.47 0.24 $1\times 4$ 88.2 70.0 E-3DSNN-S (Qiu et al., 2025a) Voxel 3.27 0.02 $1\times 4$ 91.7 78.7 E-3DSNN-L (Qiu et al., 2025a) Voxel 17.7 0.26 $1\times 4$ 91.2 80.2 Spike-driven PointFormer (Ours) Point 7.69 5.1 1 $\times$4 92.6 82.1 Spike-driven PointFormer (Ours) Point 22.1 9.4 1 $\times$4 92.1 81.7
Tab. 5 compares ANN- and SNN-based 3D backbones on ModelNet40 and ScanObjectNN, reporting input modality, parameter count, measured energy, and the temporal configuration $T{\times}D$. Among ANN models, Pointformer attains strong accuracy (92.8% on M-40 / 81.3% on Scan-O), while KPConv remains competitive on Scan-O (85.3%). Voxelized efficient baselines (E-3DANN-S) offer favorable energy profiles but slightly lower accuracy on Scan-O.
Within SNNs, prior point-based methods (e.g., P2SResLNet, SpikingPointNet, Spike PointNet) reduce energy but lag in accuracy, and voxel SNNs (E-3DSNN-S,L) strike a different tradeoff with very low energy but moderate robustness on Scan-O.
Our Spike-driven PointFormer closes the accuracy gap to leading ANN point backbones while retaining SNN efficiency. The 7.69M-parameter variant achieves 92.6% on M-40 and 82.1% on Scan-O at 5.1 mJ with a $1{\times}4$ temporal setup, surpassing Pointformer on Scan-O with substantially lower energy. The larger 22.1M variant yields similar accuracy (92.1% / 81.7%) at 9.4 mJ. These results indicate that spike-driven transformers can deliver ANN-level recognition performance on point clouds while maintaining the energy advantages characteristic of SNNs.